Initially 100milligrams of a radioactive substance was present. After 6 hours the mass has decreased by 3%. The rate of decay is proportional to the amount of the substance present at time t. Determine the half-life of the radioactive substance. 


1
Expert's answer
2021-08-26T07:19:53-0400

Given:

"m_0=100\:\rm mg"

"m=97\:\rm mg"

"t=6\:\rm hr"


The radioactive decay equation (https://en.wikipedia.org/wiki/Radioactive_decay) says

"N=N_0*2^{-t\/t_{1\/2}}"

The mass of substance is proportional to the number of atoms

"m=\frac{N}{N_A}M"

So

"m=m_0*2^{-t\/t_{1\/2}}"

Hence, the half-life of the radioactive substance

"t_{1/2}=-t/\log_2\frac{m}{m_0}\\=-6\:\rm hr/\log_2\frac{97}{100}=137\: hr=5.7 \: days"